OpenStudy (anonymous):

A student answered 18 of the 120 questions on a test incorrectly. What percent of the questions were answered correctly?

OpenStudy (lgbasallote):

how much did he answer correctly?

OpenStudy (anonymous):

Assume he answered 102 questions correctly..:)

OpenStudy (anonymous):

to find out the percentage of the questions that were answered correctly, we should find out the number of correct answer. It can be found by (120-18) the percentage of correct answer =((no. of correct answer) / (total number))x100%

OpenStudy (anonymous):

ok i did that & got 17/20

OpenStudy (anonymous):

don't forget to x100%

OpenStudy (anonymous):

i did 120-18=102 then 102/120 * 100% and got 17/20.

OpenStudy (theviper):

\[\LARGE{\frac{120-18}{120}\times100}\]

OpenStudy (theviper):

\[\LARGE{\color{blue}{\frac{102}{120}\times100}}\]

OpenStudy (theviper):

solve it??

OpenStudy (anonymous):

85?

OpenStudy (theviper):

yeah

OpenStudy (anonymous):

THX!!!

OpenStudy (theviper):

yw:)

OpenStudy (anonymous):

All ready am ;)

OpenStudy (anonymous):

lol

OpenStudy (theviper):

lol